OpenStudy (anonymous):

Solve. 5 = –(–z + 3) A. –8 B. 8 C. –2 D. 2

OpenStudy (solomonzelman):

5 = –(–z + 3) 5= x-3 2=x It is D

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Anytime!